OpenStudy (anonymous):

X=w+y+z/4 sole for y

OpenStudy (anonymous):

As stated, X=w+y+z/4 .. subtracting w and z/4 from both sides would be X-w-z/4 = y However, if you meant X=(w+y+z)/4 then you multiply both sides by 4 and then subtract w and z giving you 4X-w-z=y